OpenStudy (anonymous):

finding x, y, z? i really really don't understand this one. none of the answers i get make sense :/

OpenStudy (kainui):

Sure, I can help you out.

OpenStudy (anonymous):

angle one equals 18x, angle two is x + y and angle three is 30 z

OpenStudy (kainui):

Well what can you say about those angles? Can you use that information to make any equations? I'll help walk you through it.

OpenStudy (anonymous):

i know that angle one and angle two equal 180 and angle two and three equal 180

OpenStudy (kainui):

Alright cool, so write out what you said but as an equation instead of a sentence and we'll try solving them.

OpenStudy (anonymous):

18x + x + y = 180 and x + y + 30z = 180

OpenStudy (kainui):

Alright now it looks like you need a 3rd equation to solve this one. Do you know what vertical angles are?

OpenStudy (anonymous):

i do

OpenStudy (kainui):

Can you apply that here and make another equation with it?

OpenStudy (anonymous):

18x + 30z = 180

OpenStudy (kainui):

Almost, those are the angles involved, but they don't equal 180 degrees. These are vertical angles: 

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

18x = 30z

OpenStudy (kainui):

Awesome, now I believe you have enough information to substitute in your equations to solve for all your variables, yeah?

OpenStudy (anonymous):

do they all equal 90?

OpenStudy (kainui):

I'm not sure, but I do know you have enough information to solve for all of your variables now so it's not important.

OpenStudy (anonymous):

Its a rhombus.. .in rhombus diagonals are perpendicular to each other.. !!!! so all angles are 90 degrees!